OpenStudy (anonymous):

[3x+1]/16+[2x-3]/7=[x+3]/8+[3x-1]/14 Value for X please

OpenStudy (anonymous):

1/7 (2 x-3)+1/16 (3 x+1) = (x+3)/8+1/14 (3 x-1) Expand out terms on both sides: (53 x)/112-41/112 = (19 x)/56+17/56 Add (41/112-(19 x)/56) to both sides: (15 x)/112 = 75/112 Divide both sides by 15/112: x = 5

OpenStudy (anonymous):

thanks a lot